OpenStudy (meilendurcer):

A class has 27 students. Fifteen are boys and the rest are girls. What is the ratio of girls to total students?

OpenStudy (anonymous):

12:27

OpenStudy (anonymous):

total no. of students =27 no. of boys=15 no. of girls=27-15=12 required ratio=12/27=4/9

OpenStudy (anonymous):

required ratio=4/9 0r 4:9